# Baryonium Dark Matter: Events with Circle Configurations in the Atmosphere and on the Surface of Earth

### v6

@@ Abstract
This research supposes the parallel study of the manifestations of Baryonium Dark Matter (BDM) in astroparticle collisions in the atmosphere and among the footprints of the massive extraterrestrial objects that fall on the surface of Earth. These objects differ from stone asteroids and are seen as lightning bolides that are burning with a blue flame. Since BDM is a QCD state, the characteristics of BDM disintegration events can be analyzed with the same model as hadroproproduction data on colliders. It was concluded that the HE event, which was detected in the stratosphere in 1975, differs from the collision of the nucleus. The signatures of such events are 1) the circle distribution of heavy secondary hadrons with almost nothing inside the ring and 2) the presence of heavy secondary particles with a mass out of range of known hadrons. The circle will also be the signature of heavy BDM conglomerate that disintegrate in the atmosphere for a few similar lower-mass BDMs. These signatures should be observed in super-heavy BDM collapses on the Earth. The traces of such events are seen as circles of pits or shafts in the ground rocks. There are at least two pieces of evidence of such holes: cenotes on the Yucatan peninsula (Mexico) and pits in Durrington Walls (England). The vertical pits, or wells (because sometimes they are filled with water) have a depth of 5 - 50 meters. These super-heavy BDM collapses differ from the meteorite crash, because they do not form the craters. The existence of these places around the globe indicates the rare likelihood of meeting the super-massive BDM in proximity to Earth. The Tunguska meteorite could also be a BDM collapse. Finally, some suggestions are made about the emission of Baryonium Dark Matter toroidal conglomerates with the jets from gigantic active Supermassive Black Holes (SMBH).
@@ 1 Introduction into Baryonium Dark Matter
This state of baryon-antibaryon matter has been named Baryonium Dark Matter (BDM) and should have discrete levels of masses , because not every arbitrary number of baryon-antibaryon hexagons can form the torus. The tetraquarks pentaquarks,and hexaquarks includes the antiproton SJ, so can be considered as results of deconstruction of BDM torus.
In such a way, Dark Matter is an object of QCD string physics, which is responsible for baryon and antibaryon production on the wide logariphmic range of energies and masses begining from HE hadroproduction on LHC upto Ultra High Energy (UHE) matter injection with the relativistic jets from SMBH.
@@ 3 Footprints of BDM collapse events on the surface of the Earth
The supermassive Baryonium Dark Matter conglomerates could certainly reach the surface of Earth. What are we expecting to see? As it was found from the views of two locations with misterious circles, see figure 4 and figure 5: pits in Durrington Walls and cenotes on Yucatan peninsula have been organized in circles. The central areas of circles are almost empty, because, as it was seen from histograms of stratosphere event, there should be UHE light components: electromagnetic radiation, hadron showers and light components of BDM, which can not destroy the underground rock. The circle of pits is the signature of the first stage disintegration of supermassive Baryonium Dark Matter conglomerate for the number of BDMs of equal lower masses. What is important, there are no craters, as for the falls of meteorites. in our time the fall of Dark Matter is looking as lightening bolide in the sky.
@@ 6 Conclusions
The ancient pits, wells, cenotes and othes holes that are distributed as circles in the undeground rocks are visible in many places on the surface of Earth. The circles of holes are the traces of disintergration of supermassive Baryonium Dark Matter. Since QCD matter obeys to certain symmetry, the neutral BDM state can fall apart only into the identical states with discrete number of equal lower masses, which have to be organized in the circles. The Baryonium Dark Matter objects consist of ordinary baryon-antibaryon matter and may possess the value of mass in the range from nucleon mass to the mass of order Black Hole. Such matter can be built only under very high pressure in Super Massive Black Holes. We observe only heaviest parts of original BDM, because the light electromagnetic components are not leave the footprints. So these circles cannot be results of meteorit falls because of the absence of central crater. The first sign of heavy Baryonium Dark Matter with the mass above the known hadron masses was detected in 1975th in stratosphere experiment. Stratosphere event is real signature of BDM disintegration, because nothing can explain the circles in rapidity distribution at the region of intermediate rapidity and the scarce multiplicity of produced protons in forward area. Curcles are the result of disintegration of heavy mass BDM into the identical BDMs with lower masses. The likelihood exists to meet such objects in proximity of Earth. The stages of BDM disintegration in the atmosphere are similar to the BDM emission procedure with the relativistic jet: injection of BDM in jet from SMBH, the splitting of supermassive BDM into lower-mass states, and the acceleration of secondaries due to the stored energy of SJ connection. As soon as rare super- massive BDM conglomerate reaches the atmosphere of Earth, it split apart, originate hadronic shower in atmosphere, and the resulting heavy BDM debrises can punch the pits in the ground. The giant-BDM scenario, when dinosaurs went extinct, has not been realized again during few tens of million years. May be it cannot happen ever because the environment in Galaxy has been changed (more stars and dust don’t allow to penetrate through up to our solar system). On the planets without atmosphere, the fall of super-massive BDM would cause the mars quakes.

### v7

@@ Abstract
This research supposes the parallel study of the manifestations of Baryonium Dark Matter (BDM) in astroparticle collisions in the atmosphere and among the footprints of the massive extraterrestrial objects that fall on the surface of Earth. These objects differ from stone asteroids and are seen as lightning bolides that are burning with a blue flame. Since BDM is a QCD state, the characteristics of BDM disintegration events can be analyzed with the same model as hadroproproduction data on colliders. It was concluded that the HE event, which was detected in the stratosphere in 1975, differs from the collision of the nucleus. The signatures of such events are 1) the circle distribution of heavy secondary hadrons with almost nothing inside the ring and 2) the presence of heavy secondary particles with a mass out of range of known hadrons. The circle will also be the signature of heavy BDM conglomerate that disintegrate in the atmosphere for a few similar lower-mass BDMs. These signatures should be observed in super-heavy BDM collapses on the Earth. The traces of such events are seen as circles of pits or shafts in the ground rocks. There are at least two pieces of evidence of such holes: cenotes on the Yucatan peninsula (Mexico) and pits in Durrington Walls (England). The vertical pits, or wells (because sometimes they are filled with water) have a depth of 5 - 50 meters. These super-heavy BDM collapses differ from the meteorite crash, because they do not form the craters. The existence of these places around the globe indicates the rare likelihood of meeting the super-massive BDM in proximity to Earth. The Tunguska meteorite could also be a BDM collapse. Finally, some suggestions are made about the emission of Baryonium Dark Matter toroidal conglomerates with the jets from giant active Supermassive Black Holes (SMBH).
@@ 1 Introduction into Baryonium Dark Matter
This state of baryon-antibaryon matter has been named Baryonium Dark Matter (BDM) and should have discrete levels of masses , because not every arbitrary number of baryon-antibaryon hexagons can form the torus. The tetraquarks pentaquarks,and hexaquarks includes the antiproton SJ, so can be considered as results of deconstruction of neutral BDM torus.
In such a way, Dark Matter is an object of QCD string physics, which is responsible for baryon and antibaryon production on the wide logariphmic rangegyhcvf of energies and masses begining from HE hadroproduction on LHC upto Ultra High Energy (UHE) matter][ injection with the relativistic jets from SMBH.
@@ 3 Footprints of BDM collapse events on the surface of the Earth
The supermassive Baryonium Dark Matter conglomerates could certainly reach the surface of Earth. What are we expecting to see? As it was found from the views of two locations with misterious circles, see figure 4 and figure 5: pits in Durrington Walls and cenotes on Yucatan peninsula have been organized in circles. The central areas of circles are almost empty, because, as it was seen from histograms of stratosphere event, there should be UHE light components: electromagnetic radiation, hadron showers and light components of BDM, which can not destroy the underground rock. The circle of pits is the signature of the first stage disintegration of supermassive Baryonium Dark Matter conglomerate for the number of BDMs of equal lower masses. What is important, there are no craters, as for the falls of meteorites. in our time the fall of Dark Matter is looking as a flight of lightening bolide in the sky.
@@ 6 Conclusions
The ancient pits, wells, cenotes and othes holes that are distributed as circles in the undeground rocks are visible in many places on the surface of Earth. The circles of holes are the traces of disintergration of supermassive Baryonium Dark Matter. Since QCD matter obeys to certain symmetry, the neutral BDM state can fall apart only into the identical states with discrete number of equal lower masses, which have to be organized in the circles. The Baryonium Dark Matter objects consist of ordinary baryon-antibaryon matter and may possess the value of mass in the range from nucleon mass to the mass of order Black Hole. Such matter can be built only under very high pressure in Super Massive Black Holes. We observe only heaviest parts of original BDM, because the light electromagnetic components are not leave the footprints. So these circles cannot be results of meteorit falls because of the absence of central crater. The first sign of heavy Baryonium Dark Matter with the mass above the known hadron masses was detected in 1975th in stratosphere experiment. Stratosphere event is real signature of BDM disintegration, because nothing can explain the circles in rapidity distribution at the region of intermediate rapidity and the scarce multiplicity of produced protons in forward area. Curcles are the result of disintegration of heavy mass BDM into the identical BDMs with lower masses. The likelihood exists to meet such objects in proximity of Earth. The stages of BDM disintegration in the atmosphere are similar to the BDM emission procedure with the relativistic jet: injection of BDM in jet from SMBH, the splitting of supermassive BDM into lower-mass states, and the acceleration of secondaries due to the stored energy of SJ connection. As soon as rare super- massive BDM conglomerate reaches the atmosphere of Earth, it split apart, originate hadronic shower in atmosphere, and the resulting heavy BDM debrises can punch the pits in the ground. Since the components of torus are hydrogen nuclei, they are burning in oxigen atmosphere with blue flame and bring the water to the Earth. The importent question arises: how did the oxigen atmosphere appeare at the Earth? The giant-BDM scenario, when dinosaurs went extinct, has not been realized again during few tens of million years. May be it cannot happen ever because the environment in Galaxy has been changed (more stars and dust don’t allow to penetrate through up to our solar system). On the planets without atmosphere, the fall of super-massive BDM would cause the mars quakes.